# Prompt-Driven Building Footprint Extraction in Aerial Images with Offset-Building Model

### v3

@@ V Ablation
Figure 8: The influence of prompt quality. For each figure, x axis represents the box noise in pixels added by people. In (a) and (b), we discover the quality drop of predictions mainly caused by the increase of length loss. (c) tells us a suitable larger bounding box can improve the mask quality of prediction.
@@
To tap into the potential of soft DNMS, we tested the influence of different $\alpha$, and the results are shown in Table 10. and interactively use soft DNMS to improve offset quality. In conclusion, iteratively using soft DNMS has more significant improvement compared with adjusting $\alpha$.
@@ VI-A Why does OBM perform better?
Another key reason, we think, why OBM performs better on offset prediction is because of global attention. We notice that in Table 1, OBM has a better performance on those short offset. We also discover this phenomenon in our generalization test. OBM still remains it very powerful direction awareness. Far from only using features inside the prompt area like ROI based models, transformer structures in OBM still allow other parts of features involved while inferring. The final offset was not defined by a single area, but by several related areas. In other words, although shorter offsets are more difficult to predict, OBM obtains extra information from other buildings via attention. We visualized the offset tokens in Fig. 9 to show this phenomenon.
@@ VI-B Different structures of OBM
To ensure the best structure of our model. We designed four structures of OBM in Table 8. D. D. is a double decoder to predict offsets, roof and building separately. Share q. means different heads in ROAM using a shared offset token. - ROAM means we remove all adaptive heads in ROAM, only base head remained.
We tested all kinds of potential training methods to find the best performance of mentioned model the results are as Table 8. The final row is our final OBM.
TABLE 8: Different structure of our model trained on BOANI[12]
@@ VI-C Why does DNMS work?
DNMS algorithms are designed to improve offset quality by correcting offset angle. In Table 1, easily we find shorter offsets often have poor performance while longer offsets were predicted well in terms of their direction. Naturally, we rise an idea: we can use those longer prediction to correct those shorter. DNMS algorithms will not delete subjects but only improve offset angle.
After experiments, DNMS algorithms can improve offset quality by correct the angle of offsets especially for shorter offsets. This is proved by Table 2 and Table 9.
TABLE 9: In this table, DNMS and soft DNMS correct the angle for those shorter offsets. The circumstances of OBM is similar.
@@ VI-D How to use soft DNMS for better?
DNMS once used, the angles of whole image Soft DNMS is a soft version of DNMS. We need to determine the parameter $\alpha$ in soft DNMS to get the best performance. We use the prediction of LOFT as an example to show the effect of $\alpha$ in Table 10.
TABLE 10: Different $\alpha$ of soft DNMS test on BOANI[12] by LOFT
@@
As our DNMS algorithms do not delete any offsets, it is also possible to iteratively use soft DNMS. We use the prediction of LOFT as an example to show the effect of iterative soft DNMS in Table 11.
TABLE 11: Different iterations test on soft DNMS
@@ VI-E Decoupling model’s benefits
Our OBM model has more better performance than LOFT and Cascade LOFT, which is more reliable and robust. Powerful prompt encoder makes sure that OBM can integrate almost any other models, e.g. building detection, roof semantic segmentation and so on, to provide prompts. That means in real-world applications, data producers can get better results, even using the same RPN as [12] to provide prompts. For those lost target, data annotators can also label them out in a less time-consuming way because basically we can provide a roof segmentation at least.

### v4

@@ V-C The robustness of the model to prompt errors
(a) $VL$ and $LL$ / Prompting Error
(b) $AL$ / Prompting Error
(c) Roof IoU and BIoU / Prompting Error
Figure 9: The influence of prompt quality. For each figure, $x$ axis represents the prompting error added on prompting boxes, and this random noise is measured by pixels. In (a) and (b), we discover the quality drop of predictions mainly caused by the increase of length loss. (c) tells us a suitable larger bounding box can improve the mask quality of prediction.
@@
V-D Different parameters in soft DNMS
To tap into the potential of soft DNMS, we tested the influence of different $\alpha$, and the results are shown in Table XII. and interactively use soft DNMS to improve offset quality. In conclusion, iteratively using soft DNMS has more significant improvement compared with adjusting $\alpha$.
@@ VI-A Why does OBM perform better?
Another key reason, we think, why OBM performs better on offset prediction is because of global attention. We notice that in Table I, OBM has a better performance on those short offsets. We also discover this phenomenon in our generalization test. OBM still remains it very powerful direction awareness. Far from only using features inside the prompt area like ROI based models, transformer structures in OBM still allow other parts of features involved while inferring. The final offset was not defined by a single area, but by several related areas. In other words, although shorter offsets are more difficult to predict, OBM obtains extra information from other buildings via attention. We visualized the offset tokens in Fig. 10 to show this phenomenon.
@@ VI-B Different structures of OBM
To ensure the best structure of our model. We designed four structures of OBM in Table X. D. D. is a double decoder to predict offsets, roof and building separately. Share q. means different heads in ROAM using a shared offset token.
We tested all kinds of potential training methods to find the best performance of mentioned model the results are as Table X. The final row is our final OBM.
TABLE X: Different structure of our model trained on BOANI[15]
@@ VI-C Why does DNMS work?
DNMS algorithms are designed to improve offset quality by correcting offset angle. In Table I, easily we find shorter offsets often have poor performance while longer offsets were predicted well in terms of their direction. Naturally, we rise an idea: we can use those longer prediction to correct those shorter. DNMS algorithms will not delete subjects but only improve offset angle.
After experiments, DNMS algorithms can improve offset quality by correct the angle of offsets especially for shorter offsets. This is proved by Table II and Table XI.
TABLE XI: In this table, DNMS and soft DNMS correct the angle for those shorter offsets. The circumstances of OBM are similar.
@@ VI-D How to use soft DNMS for better?
DNMS once used, the angles of offsets in the whole image will be unified compulsorily. Soft DNMS is a soft version of DNMS. We need to determine the parameter $\alpha$ in soft DNMS to get the best performance. We use the prediction of LOFT as an example to show the effect of $\alpha$ in Table XII.
TABLE XII: Different $\alpha$ of soft DNMS test on BOANI[15] by LOFT
@@
As our DNMS algorithms do not delete any offsets, it is also possible to iteratively use soft DNMS. We use the prediction of LOFT as an example to show the effect of iterative soft DNMS in Table XIII.
TABLE XIII: Different iterations test on soft DNMS
@@ VI-E Decoupling model’s benefits
Our OBM model has better performance than LOFT and Cascade LOFT, which is more reliable and robust. Powerful prompt encoder makes sure that OBM can integrate almost any other models, e.g. building detection, roof semantic segmentation and so on, to provide prompts. That means in real-world applications, data producers can get better results, even using the same RPN as [15] to provide prompts. For those lost target, data annotators can also label them out in a less time-consuming way because basically we can provide a roof segmentation at least.
@@
In the past, COCO metrics, F1score, Precision and Recall were widely used in measure footprint masks[15, 78, 17, 50]; however, these metrics today may not suitable in measuring prompt-level models. As we saw in Table III and Table V, prompt-level models easily defeat those end-to-end models. The reasons are as follows: Metrics, such as F1score, Precision and Recall, indicate models’ ability of predicting an output without any references. However, the proposal network usually perform unstable: smaller IoU threshold will give out low quality instances whose masks are insufficient to cover the related items, while bounding boxes with higher IoU threshold are rare in one predicted output. In other words, end-to-end models tend to omit some important samples or give out mistake outputs. Consequently, prompt-level models avoid aforementioned problems by directly import valid bounding boxes to activate boxes regions. In other words, metrics for end-to-end models inflect the capability of finding right and good footprints, and metrics for prompt-level models will emphasize more on whether the model can give out high quality masks for each building.